OpenStudy (anonymous):

I got (3,6) is that right? Solve the system: y = 4x + 6 and 2x = y

ganeshie8 (ganeshie8):

nope. check the signs when you add subtract

OpenStudy (anonymous):

Shoot. So its -3,-6

ganeshie8 (ganeshie8):

thats right !! you can confirm this urself, by plugging in the values in the system of equations given

OpenStudy (anonymous):

Oh thanks!! I also got stuck on another one can i share it with you?

ganeshie8 (ganeshie8):

sure.. .

OpenStudy (anonymous):

Solve the system: 3x – y = 28 and y = 14 – 2x a. (-2.8, 8.4) b. (2.8, -8.4) c. (-8.4, 2.8) d. (8.4, -2.8)

ganeshie8 (ganeshie8):

hmm where you got stuck ?

OpenStudy (anonymous):

I plugged in the y equation in and i ended getting 32? i don't think thats even close.

ganeshie8 (ganeshie8):

3x – y = 28 -----------(1) y = 14 – 2x ------------(2) substitute y value from (2) in (1) => 3x - (14-2x) = 28 3x - 14 + 2x = 28 5x = 14 x = 14/5 = 2.8

OpenStudy (anonymous):

So would that make y -8.4?

ganeshie8 (ganeshie8):

ok i made a mistake. let me redo...

OpenStudy (anonymous):

Oh oh sure

ganeshie8 (ganeshie8):

3x – y = 28 -----------(1) y = 14 – 2x ------------(2) substitute y value from (2) in (1) => 3x - (14-2x) = 28 3x - 14 + 2x = 28 5x = 42 x = 42/5 = 8.4

OpenStudy (anonymous):

So that would make y 2.8?

ganeshie8 (ganeshie8):

now plugin x = 8.4 in (2) and solve for y y = 14 - 2*8.4 = -2.8

OpenStudy (anonymous):

-2.8**

OpenStudy (anonymous):

Thanks so much!!

ganeshie8 (ganeshie8):

yep!

ganeshie8 (ganeshie8):

np... yw :)